OpenStudy (anonymous):

A snowmobile has an initial velocity of 4.7 m/s. If it accelerates at the rate of 0.73 m/s2 for 5.6s, what is the final velocity? Answer in units of m/s The Answer i got was 8.788......... now there is a second part...... If instead it accelerates at the rate of -0.51 m/s2, how long will it take to reach a complete stop? Answer in units of s

OpenStudy (anonymous):

9.275 . if at retarded at 0.51 m/sec use the equation v=u+at. v=0, u=4.7,-a=-0.51m/s2. then time taken =4.7/0.51

OpenStudy (anonymous):

V=V.+at V=4.7+(0.73*5.6) V= 8.788 m/s 0= 4.7 - 0.51t t = 9.21 sec